The Robinson family and the Sanders family each used their sprinklers last summer. The water output rate for the Robinson family's sprinkler was 25L per hour. The water output rate for the Sanders family's sprinkler was 20L per hour. The families used their sprinklers for a combined total of 55 hours, resulting in a total water output of 1250L. How long was each sprinkler used?
Answer by mananth(16946)

x hours Robinson used
y Hours Sanders used
Total memberships
1 x + 1 y = 55.00 .............1

25 x + 20 y = 1,250.00 .............2
Eliminate y
multiply (1)by -20
Multiply (2) by 1
-20 x -20 y = -1100
25 x + 20 y = 1250
Add the two equations
5 x = 150
/ 5
x = 30.000
plug value of x in (1)
1 x + 1 y = 55
30 + y = 55
y = 55-30
y = 25
y = 25
30 hours Robinson used
25 Hours Sanders used
m.ananth@hotmail.ca